OpenStudy (anonymous):

What is the sum of the arithmetic sequence 22, 13, 4 ... if there are 28 terms? 3,710 -3,094 -2,786 4,018

OpenStudy (apoorvk):

so the first term = a = 22 common difference = difference between term and its prev. term = d = 13-22 = -9 no. of terms = n = 28 Sum of AP = S then, \[S = \frac n 2 (2a + (n-1)d)\]